3 ms·
How did you do the math on this?
by dfee 9y ago
How did you do the math on this?
- Smilliam 9y agoIndependent event probability. A 1% chance of something happening means that there is a 99% chance of the event not happening. 0.99^100 = 0.37. So a 37% chance of no brain eating out of 100 trials. 1 - 0.37 = 0.63, so a 63% chance of a brain being eaten out of 100 trials.
- jessaustin 9y ago1 - (1 - .01)^100 We're finding the chance that the study will demonstrate brain eating. The hypothesis is that brain eating is rare: .01 chance of that. So any particular subject will probably not display brain eating: .99 chance of that. Assuming brain eating tests are independent, you can raise that quantity to the hundredth power to get the chance that all 100 subjects will not display brain eating. That would be a failure of the test, so subtract that chance from unity to get the test's chance of success.
- widget1321 9y agoI'm not sure where the 50% came from off the top of my head, but I think the 73% is wrong (should be 63%). My math: If there is a 1% chance of something happening in a given sample, there is a 99% chance of it not happening. So, for a set of N samples, you multiply 0.99 * 0.99 * 0.99... N times or 0.99^N to determine the chances of it not happening in any of the samples. 0.99^100 is about 0.37. 1- 0.37 = 0.63.
- deleted 9y ago[deleted]